For an infinite-source, single server system with an arrival rate of 15 customers per hour (Poisson) and service time of 2 minutes per customer (exponential), the average number waiting in line to be served is:

a. 0.1

b. 0.133

c. 0.50

d.0.250
Answer #1

c)0.50

arrival rate = 15 customer/hour =x

  service rate = 2 min per customer

y=60/2=30

The average number in line to be served  =x/y=15/30=0.5

If x is the arrival rate and y is the service rate ,then the expected number of customers in waiting line can be expressed mathematically as x/(y-x)